# Quasi Inverse of Qubit Channels for Mixed States

### v1

@@ IV.1 The Pauli Channel
with $p_{i}\geq 0$ and $\sum_{i=0}^{3}p_{i}=1$. This leads to
where
@@
The maximum value of decrease in MSTD is given by
@@
where $\lambda_{\max}$ is the largest eigenvalue of matrix $Q$ in Eq. (15). The normalized eigenstate $(x_{0},~{}\mathbf{x})^{T}$ corresponding to this largest eigenvalue will determine the quasi inverse which is given by
$V=i\sigma_{i}\quad\text{ if ~{}}p_{i}=p_{\max}\quad i\in\{1,2,3\}\,.$ (17)
In this case, the decrease in the MSTD is given by
@@
where $p_{\max}=\max(p_{1},p_{2},p_{3})$. Therefore, if $p_{\max}>p_{0}$, the quasi inverse of the Pauli channel exists. We find that if $p_{0}\leq 1/2$ and $p_{\max}\geq 1/2$, then $\Delta\overline{D^{2}}$ is maximized.
IV.2 Generalized Amplitude Damping Channel
@@
$E_{1}$ $=\sqrt{p}\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}\,,$ (19)
$E_{2}$ $=\sqrt{p}\begin{pmatrix}0&\sqrt{\gamma}\\ 0&0\end{pmatrix}\,,$
$E_{3}$ $=\sqrt{1-p}\begin{pmatrix}\sqrt{1-\gamma}&0\\ 0&1\end{pmatrix}\,,$
$E_{4}$ $=\sqrt{1-p}\begin{pmatrix}0&0\\ \sqrt{\gamma}&0\end{pmatrix}\,.$
This leads to
where
$Q=\frac{1}{2}\text{diag}\big{(}0,-\alpha(\alpha+1),-\alpha(\alpha+1),-2\alpha\big{)}$ (21)
and $\alpha^{2}={1-\gamma}$.
Clearly, the largest eigenvalue $\lambda_{\max}=-\alpha$ for $\alpha<0$ and thus the quasi inverse is
$V=i\sigma_{3}\,.$ (22)
In this case, the maximum decrease in MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{cases}0\,,&\quad\text{for }\alpha>0\,,\\ -\alpha\,,&\quad\text{for }\alpha<0\,.\end{cases}$ (23)
IV.3 Mixed Unitary Channel
@@
where $U_{i}=\exp(-i\frac{\theta}{2}\sigma_{i})=\cos(\frac{\theta}{2})I-\sin(\frac{\theta}{2}\sigma_{i})$ is a unitary operator that rotates the qubit around $x_{i}$ axis by an angle $\theta$. This leads to
where
$Q=\begin{pmatrix}0&v/2&v/2&v/2\\ v/2&q&0&0\\ v/2&0&q&0\\ v/2&0&0&q\end{pmatrix}\,,$ (26)
$v=p\sin(\theta)$, and $q=4p\sin[2](\frac{\theta}{2})-1$. For $q\geqslant 0$, the largest eigenvalue of this matrix is $\lambda_{\max}=\frac{1}{2}(q+$ $\sqrt{q^{2}+3v^{2}}$ ), with the corresponding eigenvector given by $\left(\frac{3v}{2\lambda_{\max}}\quad 1\quad 1\quad 1\right)^{T}$. This means that the quasi inverse of the channel is
@@
$\cos\phi=\frac{\sqrt{3}v}{\sqrt{3v^{2}+4\lambda_{\max}^{2}}},\quad\mathbf{n}=\frac{1}{\sqrt{3}}(x+y+z)\,.$ (28)
The decrease in MSTD is given by
@@ IV.4 Tetrahedron Channel
where $p+p^{\prime}\leq 0.5$ due to normalization of probability. This leads to:
where
$Q=\begin{pmatrix}0&0&0&0\\ 0&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1&-\frac{2p}{3}+\frac{2p^{\prime}}{3}&0\\ 0&-\frac{2p}{3}+\frac{2p^{\prime}}{3}&\frac{8p}{3}+\frac{8s}{3}-1&0\\ 0&0&0&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1\end{pmatrix}\,.$ (34)
@@
The quasi inverse for such channel is thus given by:
@@
V Conclusions
We proposed an alternative definition for quasi inverse of a channel that can easily extend the quasi inverse of channels to mixed input states. The quasi inverse based on maximizing the mean squared trace distance (MSTD) over the full Bloch ball of input state of a completely positivie and trace preserving channel was postulated to be a unitary operator. This resulted in a constrained optimization program to find the four real paramters of the inverse operator. The results for Pauli, generalized amplitude damping, mixed unitary, and the tetrahadron channels agreed with those derived based on the fidelity maximization of a channel. Therefore, this new definition provides an equivalent methodology to find the quasi inverse that is also generalizable to mixed state.

### v3

@@ V.2 The Pauli Channel
with $p_{i}\geq 0$ and $\sum_{i=0}^{3}p_{i}=1$.
The affine map for Pauli channel can be computed as
$\begin{gathered}M=\begin{pmatrix}\alpha_{23}&0&0\\ 0&\alpha_{13}&0\\ 0&0&\alpha_{12}\end{pmatrix}\quad,\quad\mathbf{c}=\begin{pmatrix}0&0&0\end{pmatrix}^{T}\,,\end{gathered}$ (33)
where $\alpha_{ij}=1-2p_{i}-2p_{j}$.
Using Eqs. (14), (17), (21), and (33) in Eq. (10) we get
$\begin{gathered}\Delta\overline{D^{2}}=\frac{2}{5}\Big{(}-p_{1}x_{0}^{2}+p_{1}x_{1}^{2}+p_{1}-p_{2}x_{0}^{2}+p_{2}x_{2}^{2}+p_{2}-\\ p_{3}x_{0}^{2}+p_{3}x_{3}^{2}+p_{3}+\frac{3}{4}x_{0}^{2}-\frac{1}{4}x_{1}^{2}-\frac{1}{4}x_{2}^{2}-\frac{1}{4}x_{3}^{2}-\frac{3}{4}\Big{)}\,.\end{gathered}$ (34)
Using the identities $x_{0}^{2}+\mathbf{x}\cdot\mathbf{x}=1$ and $\sum_{i=0}^{3}p_{i}=1$, the above expression takes the form of Eq. (26) with
@@
A detailed derivation of Eq. (35) from Eq. (34) is presented in Appendix B.
The maximum value of decrease in the MSTD is thus given by
@@
where $\lambda_{\max}$ is the largest eigenvalue of matrix $Q$ in Eq. (35). The normalized eigenstate $(x_{0},~{}\mathbf{x})^{T}=\mathbf{e}_{i}$ with $\mathbf{e}_{0}=(1,0,0,0)$, $\mathbf{e}_{1}=(0,1,0,0)$, $\mathbf{e}_{2}=(0,0,1,0)$, and $\mathbf{e}_{3}=(0,0,0,1)$ corresponding to the largest eigenvalue will determine the quasi inverse, as explained in the appendix, which is given by
$V=\sigma_{i}\quad\text{ if ~{}}p_{i}=p_{\max}\,,\quad i\in\{1,2,3\}\,,$ (37)
after using Eq. (13). In this case, the decrease in the MSTD is given by
@@
where $p_{\max}=\max(p_{1},p_{2},p_{3})$. Therefore, the quasi inverse of the Pauli channel exists if $p_{\max}>p_{0}$. We find that if $p_{0}\leq 1/2$ and $p_{\max}\geq 1/2$, then $\Delta\overline{D^{2}}$ is maximized.
V.3 Generalized Amplitude Damping Channel
@@
$E_{1}$ $=\sqrt{p}\begin{pmatrix}1&0\\ 0&\gamma\end{pmatrix}\,,$ (39)
$E_{2}$ $=\sqrt{p}\begin{pmatrix}0&\sqrt{1-\gamma^{2}}\\ 0&0\end{pmatrix}\,,$
$E_{3}$ $=\sqrt{1-p}\begin{pmatrix}\gamma&0\\ 0&1\end{pmatrix}\,,$
$E_{4}$ $=\sqrt{1-p}\begin{pmatrix}0&0\\ \sqrt{1-\gamma^{2}}&0\end{pmatrix}\,.$
The affine map for this channel can be worked out as
$M=\begin{pmatrix}\gamma&0&0\\ 0&\gamma&0\\ 0&0&\gamma^{2}\end{pmatrix},\quad\mathbf{c}=\begin{pmatrix}0\\ 0\\ (1-\gamma^{2})(2p-1)\end{pmatrix}\,.$ (40)
Note that the $p$-dependence is present only in the translational vector $\mathbf{c}$ and matrix $M$ is independent of $p$.
Using Eqs. (14) and (17) in Eq. (10), we get
$\Delta\overline{D^{2}}$ $=\frac{1}{20}\left[\text{Tr}(MM^{\dagger}-NN^{\dagger})-2\text{Tr}(M-N)\right]+\frac{1}{4}\left(\absolutevalue{\mathbf{c}}^{2}-\absolutevalue{\mathbf{u}}^{2}\right)$ (41)
$=\frac{1}{20}\left[\text{Tr}(MM^{\dagger}-NN^{\dagger})-2\text{Tr}(M-N)\right]+\frac{1}{4}\left(\absolutevalue{\mathbf{c}}^{2}-\absolutevalue{M^{i}\mathbf{c}}^{2}\right)$
[…]
$\Delta\overline{D^{2}}=\frac{1}{20}\left[\text{Tr}(MM^{\dagger}-NN^{\dagger})-2\text{Tr}(M-N)\right]\,.$ (42)
Since $M$ is indpendent of $p$, so is $N=M^{i}M$. Therefore, $\Delta\overline{D^{2}}$ is independent of $p$.
Using Eqs., (18), (21), and (40) in the above expression gives:
$\begin{gathered}\Delta\overline{D^{2}}=\frac{1}{10}\left[x_{0}^{2}(\gamma^{2}+2\gamma)-x_{1}^{2}\gamma^{2}-x_{2}^{2}\gamma^{2}+\right.\\ \left.x_{3}^{2}(\gamma^{2}-2\gamma)-\gamma^{2}-2\gamma\right]\end{gathered}\,.$ (43)
Using the identity $x_{0}^{2}+\mathbf{x}\cdot\mathbf{x}=1$, the above expression also takes the form of Eq. (26) with
$Q=\frac{1}{2}\text{diag}\big{(}0,-\gamma(\gamma+1),-\gamma(\gamma+1),-2\gamma\big{)}\,.$ (44)
When $\gamma\geq 0$, $\lambda_{\max}=0$ with corresponding eigenstate $\mathbf{e}_{0}=(1,0,0,0)$ that results in $V=I$, that is, no quasi inverse exists for positive values of $\gamma$. Clearly, the largest eigenvalue $\lambda_{\max}=-{\gamma}$ for ${\gamma}<0$ with corresponding eigenstate $\mathbf{e}_{3}=(0,0,0,1)$ that leads to $V=\sigma_{3}$ and thus
$V=\begin{cases}I\,,&\quad\text{for }\gamma\geq 0\,,\\ \sigma_{3},&\quad\text{for }\gamma<0\,,\end{cases}$ (45)
in agreement with Ref. [2]. In this case, the maximum decrease in the MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{cases}0\,,&\quad\text{for }\gamma\geq 0\,,\\ -\gamma\,,&\quad\text{for }\gamma<0\,.\end{cases}$ (46)
For example, for $p=1$ and $\gamma=0$, $Q=\mathrm{diag}(0,0,0,0)$, which means $\Delta\overline{D^{2}}=0$ for which $V=I$.
The generalized amplitude damping channel is composed of dissipation part and the phase-flip part. The quasi inverse, being a unitary map, only reverses the phase-flip part by applying $\sigma_{3}$ when $\gamma$ is negative. So, the dissipation part of the channel cannot be reversed.
V.4 Mixed Unitary Channel
@@
where $U_{i}=\exp(-i\frac{\theta}{2}\sigma_{i})=\cos(\frac{\theta}{2})I-\sin(\frac{\theta}{2})\sigma_{i}$ is a unitary operator that rotates the qubit around $x_{i}$ axis by an angle $\theta$.
The affine map for this channel is found as
$M=\begin{pmatrix}-q&-v&v\\ v&-q&-v\\ -v&v&-q\end{pmatrix}\quad,\quad\mathbf{c}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}\,,$ (48)
where $v=p\sin(\theta)$ and $q=4p\sin[2](\frac{\theta}{2})-1$. Using Eqs. (14), (17), (21), and (48) in Eq. (10) we get
$\begin{gathered}\Delta\overline{D^{2}}=\frac{2}{5}\left[\frac{q}{4}\left(-3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{3}\right)+\right.\\ \left.v(x_{0}x_{1}+x_{0}x_{2}+x_{0}x_{3})+\frac{3q}{4}\right]\,.\end{gathered}$ (49)
This allows us to write the above expression in the matrix form given by Eq. (26) with
$Q=\begin{pmatrix}0&v/2&v/2&v/2\\ v/2&q&0&0\\ v/2&0&q&0\\ v/2&0&0&q\end{pmatrix}\,.$ (50)
For $q\geqslant 0$, the largest eigenvalue of this matrix is $\lambda_{\max}=\frac{1}{2}(q+$ $\sqrt{q^{2}+3v^{2}}$ ), with the corresponding eigenstate given by $\left(\frac{3v}{2\lambda_{\max}}\quad 1\quad 1\quad 1\right)^{T}$. This means that the quasi inverse of the channel is
@@
$\cos\phi=\frac{\sqrt{3}v}{\sqrt{3v^{2}+4\lambda_{\max}^{2}}},\quad\mathbf{n}=\frac{1}{\sqrt{3}}(\hat{x}+\hat{y}+\hat{z})\,.$ (52)
The decrease in the MSTD is given by
@@ V.5 Tetrahedron Channel
where $p+p^{\prime}\leq 0.5$ due to normalization of probability.
The affine map for this channel is found as
$\begin{gathered}M=\begin{pmatrix}-\frac{8p}{3}-\frac{8p^{\prime}}{3}+1&-\frac{4p}{3}+\frac{4p^{\prime}}{3}&0\\ -\frac{4p}{3}+\frac{4p^{\prime}}{3}&-\frac{8p}{3}-\frac{8p^{\prime}}{3}+1&0\\ 0&0&-\frac{8p}{3}-\frac{8p^{\prime}}{3}+1\end{pmatrix}\,,\\ \mathbf{c}=\begin{pmatrix}0&0&0\end{pmatrix}^{T}\,.\end{gathered}$ (57)
Using Eqs. (14), (17), (21), and (57) in Eq. (10) we get
$\Delta\overline{D^{2}}=\frac{2}{5}\left[\left(\frac{8p}{3}+\frac{8p^{\prime}}{3}-1\right)(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})+\frac{4}{3}(p^{\prime}-p)x_{1}x_{2}\right]\,,$ (58)
which can be written in the matrix form given in Eq. (26) with
$Q=\begin{pmatrix}0&0&0&0\\ 0&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1&-\frac{2p}{3}+\frac{2p^{\prime}}{3}&0\\ 0&-\frac{2p}{3}+\frac{2p^{\prime}}{3}&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1&0\\ 0&0&0&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1\end{pmatrix}\,.$ (59)
@@
with corresponding eigenstates given as
$(x_{0},\mathbf{x})=\begin{cases}\frac{1}{\sqrt{2}}(0,1,1,0)&\text{ if }p\geq p^{\prime}\,,\\ \frac{1}{\sqrt{2}}(0,1,-1,0)&\text{ if }p\leq p^{\prime}\,.\end{cases}$ (61)
Therefore, the quasi inverse is
@@
VI Conclusions
We proposed an alternative definition for quasi inverse of a channel that can easily extend the mixed input states. The quasi inverse based on maximizing the mean squared trace distance (MSTD) over the full Bloch ball of input state of a completely positive and trace preserving channel was postulated to be a unitary operator. This resulted in a constrained optimization program to find the four real parameters of the inverse operator. The results for Pauli, generalized amplitude damping, mixed unitary, and the tetrahedron channels agreed with those derived based on the maximization of the average fidelity of a channel. As an additional test for the proposed method, the quasi inverse of a general unitary operator was found to be equal to the inverse of the unitary operator. Therefore, this new definition provides an equivalent methodology to find the quasi inverse that is also generalizable to mixed states.
Appendix A Maximization
To maximize $f(\mathbf{x})=\mathbf{x}^{T}Q\mathbf{x}$ subject to the constraint $\|\mathbf{x}\|^{2}=1$, we can use a Lagrange function
$\mathcal{L}(\mathbf{x},\lambda)$ $=\mathbf{x}^{T}Q\mathbf{x}-\lambda\left(\|\mathbf{x}\|^{2}-1\right)\,.$ (A.1)
Next, we need to compute partial derivatives of $\mathcal{L}$ with respect to each component of $\mathbf{x}$ and $\lambda$ and set them equal to 00 to get
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}$ $=2Q\mathbf{x}-\lambda(2\mathbf{x})=0\quad\Rightarrow\quad Q\mathbf{x}=\lambda\mathbf{x}\,,$ (A.2)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=\|\mathbf{x}\|^{2}-1=0\quad\Rightarrow\quad\|\mathbf{x}\|^{2}=1\,.$
Therefore, the problem of optimization of $f(\mathbf{x})=\mathbf{x}^{T}Q\mathbf{x}$ has been converted to an eigenvalue problem of matrix $Q$ and the solution of the optimization problem $\mathbf{x}$ and $\lambda$ must be an eigen pair of $Q$. Since $Q$ has four eigenstates $\mathbf{x}=\mathbf{v}_{i}$ with corresponding eigen values $\lambda=\lambda_{i}$, $i=1,2,3,4$. For each solution, $f(\mathbf{x})$ becomes
$f(\mathbf{x})=\mathbf{v}_{i}^{T}\left(\lambda_{i}\mathbf{v}_{i}\right)=\lambda_{i}\mathbf{v}_{i}^{T}\mathbf{v}_{i}=\lambda_{i}\,,\quad i=1,2,3,4.$ (A.3)
Thus
$\max f(\mathbf{x})=\max\left(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\right)\,,$ (A.4)
and the eigenstate $\mathbf{v}_{j}$ corresponding to $\lambda_{j}=\lambda_{\max}$ is the global maximizer of $f(\mathbf{x})$.
Appendix B A Sample Derivation
We now derive Eq. (35) from Eq. (34). Gathering the coefficients of $x_{i}^{2}$ in Eq. (34) and using the identity $\sum_{i=0}^{3}p_{i}=1$, we get:
$\Delta\overline{D^{2}}=\frac{2}{5}\left[\sum_{i=0}^{3}\left(p_{i}-\frac{1}{4}\right)x_{i}^{2}+\left(\frac{1}{4}-p_{0}\right)\right]\,,$
which can be rearranged in matrix form as
$\Delta\overline{D^{2}}$ $=\frac{2}{5}\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}\begin{pmatrix}p_{0}-\frac{1}{4}&0&0&0\\ 0&p_{1}-\frac{1}{4}&0&0\\ 0&0&p_{2}-\frac{1}{4}&0\\ 0&0&0&p_{3}-\frac{1}{4}\end{pmatrix}\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}$
$+\frac{2}{5}\left(\frac{1}{4}-p_{0}\right)\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}\mathbb{I}_{4\times 4}\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}$
$=\frac{2}{5}\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}Q\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}$
where $Q=\mathrm{diag}(0,p_{1}-p_{0},p_{2}-p_{0},p_{3}-p_{0})$ and $\mathbb{I}_{4\times 4}$ is $4\times 4$ identity matrix. A similar derivation holds for other examples as well.